# ReLoop: "Seeing Twice and Thinking Backwards" via Closed-loop Training to Mitigate Hallucinations in Multimodal understanding

### v1

@@ Abstract
While Multimodal Large Language Models (MLLMs) have achieved remarkable progress in open-ended visual question answering, they remain vulnerable to hallucinations. These are outputs that contradict or misrepresent input semantics, posing a critical challenge to the reliability and factual consistency. Existing methods often rely on external verification or post-hoc correction, lacking an internal mechanism to validate outputs directly during training. To bridge this gap, we propose ReLoop, a unified closed-loop training framework that encourages multimodal consistency for cross-modal understanding in MLLMs. ReLoop adopts a ring-shaped structure that integrates three complementary consistency feedback mechanisms, obliging MLLMs to "seeing twice and thinking backwards". Specifically, ReLoop employs the frozen Consistency Feedback Plugin (CFP), comprising semantic reconstruction, visual description, and an attention supervision module for attention alignment. These components collectively enforce semantic reversibility, visual consistency, and interpretable attention, enabling the model to correct its outputs during training. Extensive evaluations and analyses demonstrate the effectiveness of ReLoop in reducing hallucination rates across multiple benchmarks, establishing a robust method for hallucination mitigation in MLLMs. We will release our source code and data in the camera-ready version.
@@
Jianjiang Yang1thanks: Equal contribution, first author., Ziyan Huang2thanks: Equal contribution, second author., Yanshu Li3 1Department of Computer Science, University of Bristol 2School of Future Technology, South China University of Technology 3Brown University edisonyang109@gmail.com, bonnie.ziyan.huang@gmail.com, yanshu_li1@brown.edu
@@ 1 Introduction
To address this issue, we propose ReLoop, a cognitively inspired unified training framework that encourages multimodal consistency for cross-modal understanding in MLLMs. ReLoop implements a feedback-driven closed-loop supervision process, allowing the model to reassess its predictions and validate their consistency with the original input through multi-level supervision during training. Specifically, after MLLMs produce an answer from the image-question pair, Reloop enables the model to recapture input semantics and assess internal consistency via: a Consistency Feedback Plugin (CFP), comprising two frozen modules: (1) CFP-Lang reconstructs the question $\hat{Q}^{*}$ from $(A,I)$ to supervise semantic alignment, and (2) CFP-Vis generates a description ${I}^{*}$ to assess factual grounding. In parallel, an attention supervision module extracts the model’s token-to-image attention map $\mathcal{H}$ and compares it with an entropy-based pseudo-ground truth. All signals are integrated as differentiable losses in the overall optimization objective. This design encourages the model to "see twice and think backward"—first see to answer ($Q,I\rightarrow A$), see twice to reassess ($A,I\rightarrow\hat{Q}^{*},{I}^{*},\mathcal{H}$), and finally to correct ($\hat{Q}^{*},{I}^{*},\mathcal{H}$ $\approx Q,I,\mathcal{H}_{\text{pseudo}}$).
@@
Figure 2: Seeing Twice and Thinking Backwards: ReLooping Hallucination Suppression in Multimodal Language Models. This diagram aligns human cognitive phases (left) with model modules (right) in a closed-loop process. The main model $M$ produces an answer which is then introspected via CFP-Lang (language reconstruction), CFP-Vis (visual description), and internal cross-attention maps. Semantic aggregation, CLIP similarity, and entropy-based soft masks produce feedback losses that are summed and back-propagated to update $M$ and the semantic aggregator $S$.
@@ 2 Related Work
Multimodal LLMs frequently produce hallucinations—responses conflicting with visual inputs, such as inventing entities or misaligning semantics (Li et al., 2023b). Recent mitigation efforts combine post-hoc correction and architectural refinement. Retrieval-augmented methods like (Mala et al., 2025) grounds outputs in external knowledge via hybrid retrievers, while (Ayala and Béchard, 2024) reduces hallucinations in structured outputs. Architectural solutions such as OPERA (Huang et al., 2024) penalize over-trust during decoding, and preference-aligned training like TPO (Gu and Wang, 2025) enhances vision grounding. Post-generation frameworks, including Woodpecker (Yin et al., 2023), further improve factuality through structured verification.
@@
Human cognition leverages bidirectional reasoning to validate hypotheses—a principle termed "cognitive reversibility" (Johnson-Laird, 1983). Recent works explore this idea through decoding-time strategies: Self-RAG (Asai et al., 2023) integrates retrieval-augmented generation with self-reflection, enabling models to critique and refine their outputs iteratively, while DeepSeekMath employs Group Relative Policy Optimization (GRPO) (Shao et al., 2024), enhancing mathematical reasoning by optimizing policy decisions based on group sampling strategies. Similarly, back-translation methods (Sennrich et al., 2016) enforce answer-question consistency through round-trip translation. Training-time solutions remain limited: CycleConsistency (Pang and Wang, 2020) aligns forward-backward pathways via joint training but struggles with error accumulation in open-domain settings.
@@
Ensuring cross-modal consistency is vital for mitigating hallucinations in multimodal large language models (MLLMs). Recent methods enhance visual-text alignment to reduce semantic drift. Visual Contrastive Decoding (VCD) (Leng et al., 2024) contrasts outputs from original and perturbed images to promote grounding and reduce unimodal bias. Hallucination-Augmented Contrastive Learning (HACL) (Jiang et al., 2024) treats hallucinated captions as hard negatives to improve alignment. EAGLE (Villa et al., 2025) further refines visual encoders post-pretraining, yielding better grounding and fewer hallucinations.
@@ 3.2 Consistency Signals
4 ReLoop Framework: Reflect, Recapturing, and Optimize through a Closed-Loop Process
We introduce ReLoop, a unified training framework aimed at reducing hallucinations in MLLMs for open-ended VQA answering. As illustrated in Figure 2, the framework incorporates three complementary consistency feedback mechanisms: semantic reconstruction, visual description, and attention alignment to supervise the model toward producing answers faithful to both the question and the image.
These feedback signals are instantiated through a frozen Consistency Feedback Plugin (CFP): semantic reconstruction (CFP-Lang) and visual description (CFP-Vis), and attention supervision from the model itself. The CFP module is broadly compatible with a range of encoder-decoder or decoder-only MLLMs. During inference (First See → Answer), the model receives a question-image pair and produces an initial answer. The training process then begins with Reflect → Second See → Correct: the model examines its output through structured consistency feedback. Specifically, it "introspectively" asks:
  • "Did I understand the $Q$?" (→ semantic reconstruction)
@@ 4 ReLoop Framework: Reflect, Recapturing, and Optimize through a Closed-Loop Process
  • "Did I focus the right region?" (→ attention alignment)
@@ 4.1 A Closed-loop Training
    First See: The main model $M$ takes the image $I$ and question $Q$ as input to produce an initial answer $A$.
@@
    Second See: The reconstructions are compared against the original inputs to compute consistency losses, capturing discrepancies in semantics, visual grounding, and attention focus.
@@ 4.2 Re: Recapturing Details for Consistency Supervision
This stage corresponds to the training-time processes of "Reflect" and "Second See", where the model reassesses its answers to recapture overlooked semantic and visual details. Three feedback pathways modules examine whether the model understood the question, correctly grounded its answer in the image, and attended to salient regions.
@@ 4.2.3 Attention Supervision via Heatmap Consistency
To enhance interpretability and mitigate hallucinations arising from inattentive or unstable decoding, we explicitly supervise the model’s token-level cross-attention patterns. From the decoder of the main model $M$, we extract attention maps $\mathcal{H}$, which indicate the spatial focus during answer generation. We construct a soft pseudo-ground-truth heatmap $\mathcal{H}_{\text{pseudo}}$ using entropy-based masking. This method preserves uncertainty information and avoids brittle hard labels. As illustrated in Figure 3, well-grounded tokens (e.g., dog) yield concentrated heatmaps aligned with visual evidence, while hallucinated tokens (e.g., playing) produce offset patterns.
Type Module Signal Type Baseline ReLoop $\Delta$Mean Baseline Hallu. ReLoop Hallu. $\Delta$Rate
Object Visual CLIP($I$, ${I}^{*}$) 28.02 $\pm$ 3.10 29.46 $\pm$ 3.27 $\uparrow$1.44 24.5% 10.3% $\downarrow$14.2%
@@
Attention $\mathcal{H}$) 1.31 $\pm$ 0.40 1.28 $\pm$ 0.45 $\downarrow$0.03
Attribute Visual CLIP($I$, ${I}^{*}$) 26.59 $\pm$ 3.31 26.81 $\pm$ 3.41 $\uparrow$0.22 7.3% 4.0% $\downarrow$3.3%
@@
Attention $\mathcal{H}$) 1.36 $\pm$ 0.46 1.32 $\pm$ 0.52 $\downarrow$0.04
Relation Visual CLIP($I$, ${I}^{*}$) 27.22 $\pm$ 3.26 28.01 $\pm$ 3.38 $\uparrow$0.79 13.2% 7.6% $\downarrow$5.6%
@@
Attention $\mathcal{H}$) 1.39 $\pm$ 0.43 1.34 $\pm$ 0.50 $\downarrow$0.05
Event Visual CLIP($I$, ${I}^{*}$) 26.63 $\pm$ 3.08 26.94 $\pm$ 3.37 $\uparrow$0.31 10.4% 5.2% $\downarrow$5.2%
@@
Attention $\mathcal{H}$) 1.33 $\pm$ 0.42 1.51 $\pm$ 0.55 $\uparrow$0.18
@@ 5 Experimental Setup
Figure 4: Type-wise hallucination rates (%) for baseline (MiniGPT-4) and ReLoop models.
Evaluation Benchmarks and Metrics. We evaluate ReLoop on a broad range of hallucination and multimodal understanding benchmarks, including POPE (Li et al., 2023b), CHAIR (Rohrbach et al., 2018), AMBER (Wang et al., 2023), MMHal-B (Sun et al., 2023), HallusionBench (Guan et al., 2024), Faith/FaithS (Jing et al., 2024), and MME (Fu et al., 2023). Full definitions are in Appendix A.3 A.4.
@@
Baselines. We use MiniGPT-4 as the baseline model in Experiment 6.1 and compare against LLaVA-1.5 variants trained with LLaVA-RLHF Sun et al. (2023), HA-DPO Zhao et al. (2023) , and POVID Zhou et al. (2024). All baselines share the same backbone and training setup for a fair comparison. Implementation details are provided in Appendix A.5.
@@
Table 2: Performance comparison of various LVLMs with and without ReLoop. Hallucination is measured by POPE, CHAIRs, and CHAIRi, cross-modal faithfulness is evaluated using F1, Faith, and FaithS.↓ indicates lower is better; ↑ indicates higher is better.
@@ 6 Results and Analysis
We first aim to pinpoint internal representation deficiencies that drive hallucination behaviors across different hallucination types. We analyze consistency signal deviations produced by ReLoop’s frozen supervision modules, with hallucinated versus non-hallucinated samples. Responding: "Did I understand the question?” (language, via BERTScore); "Did I say it right?” (visual, via CLIP similarity); "Did I focus the right region?” (attention, via entropy).
@@
Table 4: Performance comparison of ReLoop with various alignment-enhancing baselines for LLaVA-1.5 on metrics measuring hallucination suppression and cross-modal faithfulness. Best scores are in bold and the second are underlined.
@@ 6.2 Effects of Structured Feedback in ReLoop
Hallucination Suppression. ReLoop significantly reduces references to non-existent entities. MiniGPT-4’s CHAIRs/CHAIRi drop by 20.8%/9.7% (POPE +1.9%), InstructBLIP’s by 22.7%/15.0%. LLaVA also benefits, showing a 24.9% and 33.1% drop in the same metrics. These reductions confirm that ReLoop effectively mitigates object-level and spatial hallucinations by enhancing visual consistency.
Cross-modal Faithfulness. ReLoop also enhances cross-modal faithfulness. For MiniGPT-4 and LLaVA, F1 scores improve by 10.6% and 6.8%, respectively, while Faith/FaithS rising by 4.1%, 4.9%, and 3.4% in MiniGPT-4, InstructBLIP, and ShareGPT4V. These gains suggest that the model not only grounds responses more accurately in the image but also maintains semantic alignment with the question intent.

### v2

@@ Abstract
While Multimodal Large Language Models (MLLMs) have achieved remarkable progress in open-ended visual question answering, they remain vulnerable to hallucinations. These are outputs that contradict or misrepresent input semantics, posing a critical challenge to the reliability and factual consistency. Existing methods often rely on external verification or post-hoc correction, lacking an internal mechanism to validate outputs directly during training. To bridge this gap, we propose ReLoop, a unified closed-loop training framework that encourages multimodal consistency for cross-modal understanding in MLLMs. ReLoop adopts a ring-shaped structure that integrates three complementary consistency feedback mechanisms, obliging MLLMs to "seeing twice and thinking backwards". Specifically, ReLoop employs the frozen Consistency Feedback Plugin (CFP), comprising semantic reconstruction and visual description modules, along with an attention supervision module for attention alignment. These components collectively enforce semantic reversibility, visual consistency, and interpretable attention, enabling the model to correct its outputs during training. Extensive evaluations and analyses demonstrate the effectiveness of ReLoop in reducing hallucination rates across multiple benchmarks, establishing a robust method for hallucination mitigation in MLLMs. The code is available at: https://github.com/ZiyanHuang11/Reloop-hallucinations.
@@
Jianjiang Yang1thanks: Equal contribution, Yanshu Li211footnotemark: 1, Ziyan Huang311footnotemark: 1 1University of Bristol, 2Brown University, 3South China University of Technology edisonyang109@gmail.com, yanshu_li1@brown.edu, bonnie.ziyan.huang@gmail.com
@@ 1 Introduction
To address this issue, we propose ReLoop, a cognitively inspired unified training framework that encourages multimodal consistency for cross-modal understanding in MLLMs. ReLoop implements a feedback-driven closed-loop supervision process, allowing the model to reassess its predictions and validate their consistency with the original input through multi-level supervision during training. Specifically, after MLLMs produce an answer from the image-question pair, Reloop enables the model to recapture input semantics and assess internal consistency via: a Consistency Feedback Plugin (CFP), comprising two frozen modules: (1) CFP-Lang reconstructs the question $\hat{Q}^{*}$ from $(A,I)$ to supervise semantic alignment, and (2) CFP-Vis generates a description ${I}^{*}$ to assess factual grounding. In parallel, an attention supervision module extracts the model’s token-to-image attention map $\mathcal{H}$ and compares it with an entropy-based pseudo-ground truth. All signals are integrated as differentiable losses in the overall optimization objective. This design encourages the model to "see twice and think backward"—first look to answer ($Q,I\rightarrow A$), look twice to reassess ($A,I\rightarrow\hat{Q}^{*},{I}^{*},\mathcal{H}$), and finally to correct ($\hat{Q}^{*},{I}^{*},\mathcal{H}$ $\approx Q,I,\mathcal{H}_{\text{pseudo}}$).
@@
Figure 2: Seeing Twice and Thinking Backwards: ReLooping Hallucination Suppression in Multimodal Language Models. This diagram aligns human cognitive phases (left) with model modules (right) in a closed-loop process. The main model $M$ produces an answer which is then introspected via CFP‑Lang (language reconstruction), CFP‑Vis (visual description), and internal cross‑attention maps. Semantic aggregation, CLIP similarity, and entropy‑based soft masks produce feedback losses that are summed and back‑propagated to update $M$ and the semantic aggregator $S$.
@@ 2 Related Work
Multimodal LLMs frequently produce hallucinations: responses conflicting with visual inputs, such as inventing entities or misaligning semantics (Li et al., 2023b). Recent mitigation efforts combine post-hoc correction and architectural refinement. Retrieval-augmented methods like (Mala et al., 2025) grounds outputs in external knowledge via hybrid retrievers, while (Ayala and Béchard, 2024) reduces hallucinations in structured outputs. Architectural solutions such as OPERA (Huang et al., 2024) penalize over-trust during decoding, and preference-aligned training like TPO (Gu and Wang, 2025) enhances vision grounding. Post-generation verification (Woodpecker (Yin et al., 2023)) and decoding-time personalization (PAD (Chen et al., 2024)) complement training-time alignment; concurrent work improves multimodal ICL efficiency and task control (Li et al., 2025a, b) and instruction-tuned dialog grounding (Luo et al., 2024b). Beyond images, video hallucination is diagnosed via fine-grained spatio–temporal grounding (Luo et al., 2025).
@@
Human cognition leverages bidirectional reasoning to validate hypotheses: a principle termed "cognitive reversibility" (Johnson-Laird, 1983). Recent works explore this idea through decoding-time strategies: Self-RAG (Asai et al., 2023) integrates retrieval-augmented generation with self-reflection, enabling models to critique and refine their outputs iteratively, while DeepSeekMath employs Group Relative Policy Optimization (GRPO) (Shao et al., 2024), enhancing mathematical reasoning by optimizing policy decisions based on group sampling strategies. Similarly, back-translation methods (Sennrich et al., 2016) enforce answer-question consistency through round-trip translation.
@@
Ensuring cross-modal consistency is vital for mitigating hallucinations in MLLMs. Recent methods enhance visual-text alignment to reduce semantic drift. VCD (Leng et al., 2024) contrasts outputs from original and perturbed images to promote grounding and reduce unimodal bias. HACL (Jiang et al., 2024) treats hallucinated captions as hard negatives to improve alignment. EAGLE (Villa et al., 2025) further refines visual encoders post-pretraining, yielding better grounding and fewer hallucinations. Broader benchmark efforts catalog the MLLM evaluation landscape (Li et al., 2024), including conversational aspect-based sentiment settings (Luo et al., 2024a).
@@ 3.2 Consistency Signals
4 ReLoop Framework: Reflect, Recapture, and Optimize through a Closed-Loop Process
We introduce ReLoop, a unified training framework aimed at reducing hallucinations in MLLMs for open-ended VQA answering. As illustrated in Figure 2, the framework incorporates three complementary consistency feedback mechanisms: semantic reconestruction, visual description, and attention alignment to supervise the model toward producing answers faithful to both the question and the image.
These feedback signals are instantiated through a frozen Consistency Feedback Plugin (CFP): semantic reconstruction (CFP-Lang) and visual description (CFP-Vis), and attention supervision from the model itself. The CFP module is broadly compatible with a range of encoder-decoder or decoder-only MLLMs. During inference (First Look → Answer), the model receives a question-image pair and produces an initial answer. The training process then begins with Reflect → Second Look → Correct: the model examines its output through structured consistency feedback. Specifically, it "introspectively" asks:
  • "Did I understand the QQ?" (→ semantic reconstruction)
@@ 4 ReLoop Framework: Reflect, Recapture, and Optimize through a Closed-Loop Process
  • "Did I focus on the right region?" (→ attention alignment)
@@ 4.1 A Closed-loop Training
    First Look: The main model $M$ takes the image $I$ and question $Q$ as input to produce an initial answer $A$.
@@
    Second Look: The reconstructions are compared against the original inputs to compute consistency losses, capturing discrepancies in semantics, visual grounding, and attention focus.
@@ 4.2 Re: Recapturing Details for Consistency Supervision
This stage corresponds to the training-time processes of "Reflect" and "Second Look", where the model reassesses its answers to recapture overlooked semantic and visual details. Three feedback pathways modules examine whether the model understood the question, correctly grounded its answer in the image, and attended to salient regions.
@@ 4.2.3 Attention Supervision via Heatmap Consistency
To enhance interpretability and mitigate hallucinations arising from inattentive or unstable decoding, we explicitly supervise the model’s token-level cross-attention patterns. From the decoder of the main model $M$, we extract attention maps $\mathcal{H}$, which indicate the spatial focus during answer generation. We construct a soft pseudo-ground-truth heatmap $\mathcal{H}_{\text{pseudo}}$ using entropy-based masking (Detailed explanation can be found in Appendix C). This method preserves uncertainty information and avoids brittle hard labels. As illustrated in Figure 3, well-grounded tokens (e.g., dog) yield concentrated heatmaps aligned with visual evidence, while hallucinated tokens (e.g., playing) produce offset patterns.
[…]
Object Visual CLIP(II, ${I}^{*}$) 28.02 $\pm$ 3.10 29.46 $\pm$ 3.27 $\uparrow$1.44 24.5% 10.3% $\downarrow$14.2%
@@
Attention Entropy($\mathcal{H}$) 1.31 $\pm$ 0.40 1.28 $\pm$ 0.45 $\downarrow$0.03
Attribute Visual CLIP(II, ${I}^{*}$) 26.59 $\pm$ 3.31 26.81 $\pm$ 3.41 $\uparrow$0.22 7.3% 4.0% $\downarrow$3.3%
@@
Attention Entropy($\mathcal{H}$) 1.36 $\pm$ 0.46 1.32 $\pm$ 0.52 $\downarrow$0.04
Relation Visual CLIP(II, ${I}^{*}$) 27.22 $\pm$ 3.26 28.01 $\pm$ 3.38 $\uparrow$0.79 13.2% 7.6% $\downarrow$5.6%
@@
Attention Entropy($\mathcal{H}$) 1.39 $\pm$ 0.43 1.34 $\pm$ 0.50 $\downarrow$0.05
Event Visual CLIP(II, ${I}^{*}$) 26.63 $\pm$ 3.08 26.94 $\pm$ 3.37 $\uparrow$0.31 10.4% 5.2% $\downarrow$5.2%
@@
Attention Entropy($\mathcal{H}$) 1.33 $\pm$ 0.42 1.51 $\pm$ 0.55 $\uparrow$0.18
@@ 5 Experimental Setup
Figure 4: Type‑wise hallucination rates (%) for baseline (MiniGPT-4) and ReLoop models.
Evaluation Benchmarks and Metrics. We evaluate ReLoop on a broad range of hallucination and multimodal understanding benchmarks, including POPE (Li et al., 2023b), CHAIR (Rohrbach et al., 2018), AMBER (Wang et al., 2023), MMHal-B (Sun et al., 2023), HallusionBench (Guan et al., 2024), Faith/FaithS (Jing et al., 2024), and MME (Fu et al., 2023). Full definitions can be found in Appendix A.3 A.4.
@@
Baselines. We use MiniGPT-4 as the baseline model in Experiment 6.1 and compare against LLaVA-1.5 variants trained with LLaVA-RLHF Sun et al. (2023), HA-DPO Zhao et al. (2023), POVID Zhou et al. (2024), and Visual Contrastive Decoding (VCD) Leng et al. (2023). For robustness analyses, we adopt LLaVA-1.5 + ReLoop as the canonical setting and report stress tests under noisy external supervision and nonsensical answers (Table 5). Unless otherwise specified, all baselines share the same backbone, data, and training protocol for a fair comparison. Implementation details are provided in Appendix A.5.
@@
Table 2: Performance comparison of various LVLMs with and without ReLoop. Hallucination is measured by POPE, CHAIRs, and CHAIRi; cross-modal faithfulness is evaluated using F1, Faith, and FaithS. ↓ indicates lower is better; ↑ indicates higher is better.
@@ 6 Results and Analysis
We first aim to pinpoint internal representation deficiencies that drive hallucination behaviors across different hallucination types. We analyze consistency signal deviations produced by ReLoop’s frozen supervision modules, with hallucinated versus non-hallucinated samples. Responding: "Did I understand the question?” (language, via BERTScore); "Did I say it right?” (visual, via CLIP similarity); "Did I focus on the right region?” (attention, via entropy).
@@
Table 4: Performance comparison of ReLoop with alignment-enhancing baselines for LLaVA-1.5 on hallucination suppression and cross-modal faithfulness. Best scores are in bold and the second best are underlined.
@@ 6.2 Effects of Structured Feedback in ReLoop
[…]
Cross-modal Faithfulness. ReLoop also enhances cross-modal faithfulness. F1 increases on MiniGPT-4 (+10.6%), LLaVA-1.5 (+6.8%), LLaVA-1.6 ($\sim$5.6%), and Qwen-VL-2.5($\sim$4.9%); InstructBLIP maintains comparable F1 while gaining on faith metrics. FaithS improves for MiniGPT-4 (+2.8), InstructBLIP (+3.4), Qwen-VL-2.5 (+1.3), and ShareGPT4V (+1.2), and remains near-parity on LLaVA-1.6. These gains suggest that the model not only grounds responses more accurately in the image but also maintains semantic alignment with the question intent.
6.3 Robustness Under Noisy Supervision
We next ask whether closed-loop training remains stable when external supervision is imperfect or when the initial answer signal is degenerate. We stress-test ReLoop under (i) noisy teacher feedback on the visual channel and (ii) nonsensical answers that could mislead the loop.
[…]
LLaVA-1.5 + ReLoop (clean) 87.9 42.0 19.5 52/41/9
+ Teacher noise (15%) 86.8 42.6 20.2 0.11-0.11
+ Answer noise (15%) 87.2 43.1 20.4 35/43/22 0.06-0.06
Table 5: Robustness under noisy supervision. Teacher-side visual description corruption (15%) and answer-side nonsense injection (15%) have limited impact on core hallucination metrics. CLIP similarity drops under teacher noise, whereas ACW re-allocates per-sample weights under answer noise (smaller high-confidence mass).
[…]
Natural noise robustness via fourfold filtering. ReLoop’s robustness emerges from a fourfold filter: three orthogonal supervision signals (language via BERTScore, visual via CLIP, attention via entropy-aware $\mathcal{H}_{\text{pseudo}}$) plus ACW’s discrete gating $\gamma\!\in\!\{1,0.1,0.01\}$ (Sec. 4.2). Noisy teacher feedback is first cross-validated across modalities, then attenuated by ACW, and finally diluted in the multi-loss objective (Sec. 4.3), so that biased cues have limited influence on the update direction. Empirically, the stability of POPE/CHAIR under teacher-side corruption and the expected shift in the $\gamma$ distribution under answer nonsense together indicate that the closed loop self-regularizes and converges stably despite imperfect teachers.